a. A certain battery has a 12.0 V EMF and an internal resistance of 0.005 Ω. Find the current, terminal voltage, and power dissipated by a load with 1.000 Ω.




b. How much current moves through your fingers (resistance: 1200 ohms) if you touch them to the terminals of a 6 V battery?




with solution, thank you!

1
Expert's answer
2022-04-19T15:27:33-0400

a. The current is (see https://courses.lumenlearning.com/physics/chapter/21-2-electromotive-force-terminal-voltage/):


"I = \dfrac{emf}{R_{load} +r}"

where "emf = 12V, R_{load}=1.000\Omega, r=0.005\Omega". Thus, obtain:



"I = \dfrac{12V}{1\Omega +0.005\Omega} \approx 11.9A"

The terminal voltage:



"V = emf-Ir = 12V-11.9A\cdot 0.005\Omega \approx 11.9V"

The power:


"P = I^2R_{load} = (11.9A)^2\cdot 1\Omega \approx 143W"



b. According to the Ohm's law:


"I = \dfrac{V}{R} = \dfrac{6V}{1200\Omega} = 0.005A"